III. Tutorial question. This question consists of two independent parts, A and B. A. Consider a plane wave of monochromatic light, which is shone on Pattern on screen two identical narrow slits (the widths of the slits are much less than P Q R S the wavelength of the light. The slits are separated by a distance d. The interference pattern at right is observed on a distant screen. i. If point P marks the center of the screen, what is the path length-difference between the light from each slits at points Q and S? How would your answer change if point R was the center of the screen?Explain why.
Answer #1

In youngs double slit experiment, if the center fring is brigth then the path difference for the dark fring is obtained by the follwing equation

dsin(\theta) = 2(n + 1)\frac{\lambda }{2}

where n is number of the first dark fring, second dark fring and so on...

Now in the above case, path difference for the first dark fring is (n=1)

P.D = 2(1 + 1)\frac{\lambda }{2} = 2\lambda

path difference for the second dark fring is (n=2)

P.D = 2(2 + 1)\frac{\lambda }{2} = 3\lambda

Now, path difference between first and second dark fring is:

3\lambda - 2\lambda = \lambda

Answer: Path difference between Q and S is equal to the wavelength of light